Using the Keynesian consumption function, prove numberically that, as the MPC rises from 0.7 to 0.8, saving declines. SupposeYd equals $10,000 and Co equals $1,000.
Answer #1

Consumption function is given by C = 1000 + 0.7Yd or C = 1000 + 0.7*10000 = 8000. Saving = Yd - consumption or saving = 10000 - 8000 = 2000.

Now assume that mpc = 0.8. Consumption C = 1000 + 0.8*10000 = 9000. Saving = Yd - consumption or saving = 10000 - 9000 = 1000.

Hence, as MPC rises, savings should fall because increased income is now more used in consumption that as savings.